Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, vardaro, ruggi - 16:07 Wednesday 12 January 2022 (54460) Print this report
ISC shift: DARM OLTF measurements, noise injections on longitudinal DoFs

The goal of this shift was to perform noise injections on the longitudinal DoFs at carm offset zero to characterize their OLTF, and to gather data on the relation of the DARM optomechanical TF with the alignment of the SR mirror.

At the beginning of the shift a couple of lock acquisition failed due to a ~60Hz oscillation visible on MICH (Fig.1). A quick investigation showed that the phase for B2_169MHz was detuned by about 1 rad. We kept the relative DPHI channel under control while we proceeded in the lock acquisition, and changed along the whole process. The new values are:

  • @160mW: 0.5 -> 1.3 rad
  • @480mW: 0.5 -> 1.15 rad
  • @carm_null: 0.5 ->1.2 rad

By doing so the lock acquisition worked without issue for the next attempt, and at UTC:9:03 we managed to obtain a lock that lasted 2 hours and 20 minutes. Immediately after the lock we angeged the SSFS and handed-off PRCL, MICH and SRCL to their 1f error signals.

At the beginning of this lock we adjusted the setpoints SR and BS local controls to observe their effect on the fringe (Fig.2), and engaged the dithering loops on NE and NI. We managed to obtain a better dark fringe, and we proceeded with the noise injections.

DoF Time Amplitude Filter
PRCL UTC: 9:46:00+120s 4e2


DARM UTC: 9:52:30+120s 8e1 "DARM_noise"
DARM UTC: 10:05:00+120s 2e-4 "LSC_noise_MICHband"
MICH UTC: 10:00:00+120s 4e-2 "LSC_noise_MICHband"
SRCL UTC: 10:14:45+180s 8e-2 "LSC_noise_MICHband"

After these noise injections, we began the investigation of the DARM TF while tilting the SR through the alignment interface. We set the noise on DARM using its higher band filter ("DARM_noise") and amplitude 8e1, and we moved the SR with steps of 0.2 rad in TX and TY. The SR tilt is given, in the following, as a delta with respect to the position we found at the beginning of the lock.

  • SR_TX+0.2 urad, UTC: 10:26:20+90s;
  • SR_TX+0.4 urad, UTC: 10:28:00 +90s;
  • SR_TX+0.6 urad, UTC: 10:29:50+90s;
  • SR_TX+0.8 urad, UTC: 10:31:30+90s;
  • SR_TX+1.0 urad, UTC: 10:33:30+90s;
  • SR_TX+1.2 urad, UTC: 10:35:10+90s;

During these measurements we observed that the dark fringe became better and better (Fig.3). Since moving in the opposite direction with the SR alignment to the one that is beneficial to the dark fringe is generally lethal for the lock very soon, we decided for the moment not to explore SR_TX towards negative deltas, and we used SR_TX+1.2 rad as the new working point during the scan of SR_TY:

  • SR_TY+0.2 urad, UTC: 10:41:21+90s;
  • SR_TY+0.4 urad, UTC: 10:43:10+90s;
  • SR_TY+0.6 urad, UTC: 10:45:00+90s;
  • SR_TY+0.8 urad, UTC: 10:46:50+90s;
  • SR_TY+1.0 urad, UTC: 10:48:30+90s;
  • SR_TY+1.2 urad; UTC: 10:50:20+90s;
  • SR_TY -0.2 urad; UTC: 10:55:20+90s;
  • SR_TY -0.4 urad, UTC: 10:57:50+90s;

During this scan we bserved that positive deltas affected beneficially the dark fringe power, while negative ones were detrimental (Fig.4). We kept SR_TX/Y +1.2 urad as working point, and decided to inject noise again on the longitudinal DoFs to to investigate differences in their OLTF:

DoF Time Amplitude Filter
PRCL UTC: 11:14:30+120s 4e2


DARM UTC: 11:06:00+120s 8e1 "DARM_noise"
DARM UTC: 11:10:30+120s 2e-4 "LSC_noise_MICHband"

When we tried to injecto noise on MICH, we set its amplitude to be high enough to push the corrections of MICH and SRCL (Fig.5) to the saturation limit, unlocking the ITF immediately after.

The analysis of the gathered data will continue in the following days. Figure 6 and 7 show the evolution of the DARM transfer function during the SR angular scan.

We leave the ITF with the arms locked on the IR.



Images attached to this report
Search Help